A parallel-plate capacitor consists of two metal plates, each of area 0.50 m? and of separation distance 0.25 cm. What is the capacitance of the capacitor if the space between the plates is Teflon?


1
Expert's answer
2021-03-23T08:14:26-0400

We can find the capacitance of the capacitor filled with dielectric as follows:


"C=\kappa\epsilon_0\dfrac{A}{d},""C=2.1\cdot8.85\cdot10^{-12}\ \dfrac{F}{m}\cdot\dfrac{0.5\ m^2}{0.25\cdot10^{-2}\ m}=3.72\ nF."